OpenStudy (anonymous):

14x = –36; x = –3 true or false

OpenStudy (yanasidlinskiy):

Plug in -3 for x

OpenStudy (phebe):

just plug -3 ijnto the problem and solve from there

OpenStudy (solomonzelman):

What should I say? Plug in 3 for x and compare the sides...

OpenStudy (phebe):

What do you think the answer is?

OpenStudy (yanasidlinskiy):

14 (-3) = -36 -42 DOES NOT = to -36

OpenStudy (yanasidlinskiy):

Which would be...

OpenStudy (anonymous):

I was thinking false?

OpenStudy (anonymous):

You don't need to think anything, just plug in the -3 and multiply 14 by -3...

OpenStudy (anonymous):

What do you get 14*-3?

OpenStudy (yanasidlinskiy):

Yes! You got it! FDV

OpenStudy (anonymous):

Oh I see yana already did it for you lol...

OpenStudy (anonymous):

Thanks!

OpenStudy (yanasidlinskiy):

Oh...hahaha...iambatman...at lest you helped anyways!!:) And your welcome FDV!!!:)